### v2

# Signature Isolation Forest

Abstract
Functional Isolation Forest (FIF) is a recent state-of-the-art Anomaly Detection (AD) algorithm designed for functional data. It relies on a tree partition procedure where an abnormality score is computed by projecting each curve observation on a drawn dictionary through a linear inner product. Such linear inner product and the dictionary are a priori choices that highly influence the algorithm’s performances and might lead to unreliable results, particularly with complex datasets. This work addresses these challenges by introducing Signature Isolation Forest, a novel AD algorithm class leveraging the rough path theory’s signature transform. Our objective is to remove the constraints imposed by FIF through the proposition of two algorithms which specifically target the linearity of the FIF inner product and the choice of the dictionary. We provide several numerical experiments, including a real-world applications benchmark showing the relevance of our methods.
1 Introduction
The development of sophisticated anomaly detection (AD) methods plays a central role in many areas of statistical machine learning, see Chandola et al., (2009). The area of AD is growing increasingly complex, especially in large data applications such as in network science or functional data analysis.
In the classical AD multivariate scenario, observations can be represented as points in $\mathbb{R}^{d}$ and various methodologies have been introduced in the literature. A model-based approach is favored when there is knowledge about the data-generating process (Rousseeuw and Driessen,, 1999). Conversely, in numerous real-world applications where the underlying data system is unknown, nonparametric approaches are preferred (Polonik,, 1997; Schölkopf et al.,, 2001).
The area of AD is only just beginning to explore the domain of functional data. Functional data is becoming an important discipline in topological data analysis and data science, see Ramsay and Silverman, (2005); Ferraty, (2006) or Wang et al., (2016) for more details. In essence, functional data involves treating observations as entire functions, curves, or paths. In some application domains, this approach can provide richer information than just a sequence of data observation vectors, yet it comes with challenges, especially when considering the development of AD methods.
In the use of AD within a functional data context, the anomaly detection task involves identifying which functions/curves differ significantly in the data to be considered anomalous in some sense. Following Hubert et al., (2015), three types of basic anomalies can be detected: shift, shape, amplitude. These anomalies can then be isolated/transient or persistent depending on the frequency of their occurrences. Some are easier/more difficult to identify than others, e.g. an isolated anomaly in shape is much more complex than a persistent anomaly in amplitude.
At first sight, one may adopt a preliminary filtering technique. This leads to projecting the functional data onto a suitable finite-dimensional function subspace. It can be done by projecting onto a predefined basis such as Fourier or Wavelets or on a data-dependent Karhunen-Loeve basis by means of Functional Principal Component Analysis (FPCA; Dauxois et al.,, 1982; Shang,, 2014). The coefficients describing this subspace are then used as input into a specific anomaly detection (AD) algorithm designed for multivariate data.
In this approach, the a priori choice of basis to undertake the projection can significantly influence the ability to detect particular types of anomalies in the functional data accurately. This makes these approaches sensitive to such choices. In other words, the choice of a given bases for projection can artificially amplify specific patterns or make others disappear entirely, depending on their properties. Therefore, it is a choice highly sensitive to the knowledge of the underlying studied process, which is a must for such an AD method to work.
If one seeks to avoid this projection methods that take functional data to a parameter vector space to undertake the AD method, one may alternatively seek AD methods working directly in the function space. This is the case of functional data depth, which takes as input a curve and a dataset of curves and returns a score indicating how deep the curve w.r.t. the dataset is. Many functional data depths, with different formulations, have been designed so far, such as, among others, the functional Tukey depth (Fraiman and Muniz,, 2001; Claeskens et al.,, 2014), the Functional Stahel-Donoho Outlyingness (Hubert et al.,, 2015), the ACH depth Staerman et al., (2020), the band depth (López-Pintado and Romo,, 2009) or the half-region depth (López-Pintado and Romo,, 2011); see Nieto-Reyes and Battey, (2016); Gijbels and Nagy, (2017) or the Chapter 3 of Staerman, (2022) for a detailed review of functional depths.
One may also extend the classical anomaly detection algorithms designed for multivariate data such as the one-class SVM (OCSVM; Schölkopf et al.,, 2001), Local Outlier Factor (LOF; Breunig et al.,, 2000) or Isolation Forest (IF; Liu et al.,, 2008) to the functional setting, see Rossi and Villa, (2006). Recently, Staerman et al., (2019) introduced a novel algorithm, Functional Isolation Forest (FIF), which extends the popular IF to an infinite-dimensional case. It faces the challenge of adaptively (randomly) partitioning the functional space to separate trajectories in an iterative partition. Based on the computational advantages of random tree-based structures and their flexibility in choosing the split criterion related to functional properties of the data, it has been considered a powerful and promising approach for functional anomaly detection (Staerman et al.,, 2023).
Although FIF offers a robust solution for functional data, three main challenges arise when it is applied in practice. Firstly, the selection of two critical parameters involved in the splitting criterion: the functional inner product and the dictionary of time-frequency atoms for data representation. The choice of these elements significantly influences the algorithm’s performance, potentially compromising its flexibility. Secondly, the splitting criterion is formulated as a linear transformation of the dictionary function and a sample curve, potentially constrained in capturing more complex data, including non-stationary processes. Thirdly, when dealing with multivariate functions, FIF takes into account dimension dependency only linearly by using the sum of the inner product of each marginal as a multivariate inner product and may not capture complex interactions of the multivariate process. It is crucial to address these challenges to enhance the overall efficacy and adaptability of the FIF algorithm in practical applications.
In response to the above challenges, we propose a new class of algorithms, named (Kernel-) Signature Isolation Forest, by leveraging the signature technique derived from rough path theory (Lyons et al.,, 2007; Friz and Victoir,, 2010). In rough path theory, functions are usually referred to as paths to characterize their geometrical properties. The signature transform, or signature of a path, summarizes the temporal (or sequential) information of the paths. In practice, it captures the sequencing of events and the path’s visitation order to locations. Yet, it entirely disregards the parameterization of the data, providing flexibility regarding the partial observation points of the underlying functions/paths. Consider a path in a multidimensional space of dimension $d$, i.e. $\mathbf{x}\in\mathcal{F}^{d}$, where the path consists of a sequence of data points. The signature of this path is a collection of iterated integrals of the path with respect to time. By including these iterated integrals up to a certain level, the signature provides a concise and informative representation of the path, allowing for extracting meaningful features. This makes it particularly useful in analyzing sequential data, such as time series. The idea behind such a transform is to produce a feature summary of the data system, which captures the main events of the data and the order in which they happen, without recording when they occur precisely (Lyons and McLeod,, 2022). Performances of signature-based approaches appear to be promising, as they achieve state-of-the-art in some applications such as handwriting recognition (Wilson-Nunn et al.,, 2018; Yang et al., 2016b, ), action recognition (Yang et al., 2016a, ), and medical time series prediction tasks (Morrill et al.,, 2019). Surprisingly, it has been overlooked in the Machine Learning problems and has never been used in functional anomaly detection until we make this novel connection.
Our idea is to perform anomaly detection by isolating multivariate functions corresponding to sets of paths subject to the signature transform. Hence, the tree partitioning of FIF will now take place on the feature space of the signature transform. In such a way, the functional data changes are captured by embedding the studied data set into a path whose changes are then summarized by the signature method. We propose two different algorithms named Kernel Signature Isolation Forest (K-SIF) and Signature Isolation Forest (SIF). The first algorithm extends the FIF procedure to nonlinear transformations by using the signature kernel instead of the inner product of FIF. Relying on the attractive properties of the signature, see Fermanian, (2021); Lyons and McLeod, (2022) or Section 2.2, our algorithms take into account higher moments of the functions (derivative, second-order derivative, etc.) hence accommodating changes (or oscillations) of different nature. The second algorithm is a more straightforward procedure relying on what is usually referred to as ‘coordinate signature’, defined in Section 2, and being free of any sensitive parameters such as the dictionary (for FIF and K-SIF) and the inner product (for FIF). It removes the performance variability of the parameters used in FIF.
  • We introduce two functional anomaly detection algorithms (K-SIF and SIF) based on the isolation forest structure and the signature approach from rough path theory.
  • The two algorithms provide improvements in the functional anomaly detection community by extending Functional Isolation Forest in two directions: the first one considers non-linear properties of the underlying data, hence providing a more suitable procedure for a more challenging dataset; the second one is an entirely data-driven technique free from any a priori choice due to a dictionary selection constraining the functional AD procedure to identify only specific types of patterns.
  • After studying the behavior of our class of algorithm regarding their parameters, we highlight the benefits of using (K-)SIF over FIF. On a competitive benchmark, we show that K-SIF, which extends the FIF procedure through the signature kernel, consistently showed the results of FIF on real-world datasets. Furthermore, we show that SIF achieves state-of-the-art performances while being more consistent than FIF.
2 Background & Preliminaries
Consider the functional random variable (r.v.) $\mathbf{X}$, taking values in the functional space $\mathcal{F}^{d}(I)$ of real valued functions (e.g. the Kolmogorov space $\mathcal{C}(I)$ or the Hilbert space $L_{2}(I)$) where $\textit{I}\subset\mathbb{R}$ represents an interval, such that
$\begin{split}\mathbf{X}:\Omega&\longrightarrow\mathcal{F}^{d}(\textit{I})\\ \omega&\longmapsto\mathbf{X}(\omega)=(X_{t}(\omega))_{t\in\textit{I}}.\end{split}$
[…]
In practice, only a finite-dimensional marginal, $(X_{t_{1}},X_{t_{2}},\dots,X_{t_{p}})$ with $t_{1}<\dots<t_{p}$ and $(t_{1},\dots,t_{p})\in[0,1]^{p}$, can be observed. To consider a function, rather than a set of discrete values, approximation or interpolation procedures are usually used and combined with a preliminary smoothing step when the observations are noisy. Denote by $\mathcal{X}$ a reconstructed dataset of interpolated curves/functions $\mathcal{X}=\{\mathbf{x}_{1},\ldots,\mathbf{x}_{n}\}$ from the set of observations $\{x_{i,t_{1}},x_{i,t_{2}},\dots,x_{i,t_{p}}\}_{i=1}^{n}$.The problem of functional anomaly detection can be formulated as learning a score function $s:\mathcal{F}^{d}\rightarrow\mathbb{R}$ from $\mathcal{X}$ reflecting the degree of abnormality of any element of an infinite dimensional space $\mathcal{F}^{d}$ w.r.t. $\mathcal{X}$.
2.1 Functional Isolation Forest
Consider $\mathcal{H}$ the functional Hilbert space equipped with a inner product $\langle.,.\rangle_{\mathcal{H}}$ such that any ${\bf x}\in\mathcal{H}$ is a real function defined on $[0,1]$. A Functional Isolation Forest is created through an assembly of functional isolation trees (F-itrees). Each F-itree is constructed via a series of random splits from a subsample (of size $m$) of $\mathcal{X}_{n}$. The abnormality score for an observation $\mathbf{x}$ is then determined as a monotonically decreasing transformation of $\mathbf{x}$’s average depth across the trees. The core concept lies in the randomness of the splits, where an observation markedly different from others is more likely to be isolated from $\mathcal{X}_{n}$, appearing at shallower levels in the F-itrees. The F-itrees are built based on a predetermined dictionary $\mathcal{D}\subset\mathcal{H}$, encompassing both deterministic and/or stochastic functions capturing pertinent data properties, which may also be a subset of $\mathcal{X}_{n}$. Before each random univariate split, all node observations are projected onto a line defined by a randomly selected element from the dictionary $\mathcal{D}$. The selection of a suitable dictionary plays a pivotal role in shaping the FIF score construction. The projection criterion at each node of each F-itree is defined as:
$\langle\mathbf{x},\mathbf{d}\rangle_{\mathcal{H}}=\alpha\times\dfrac{\langle\mathbf{x},\mathbf{d}\rangle_{{}_{L_{2}}}}{||\mathbf{x}||||\mathbf{d}||}+(1-\alpha)\times\dfrac{\langle\mathbf{x^{\prime}},\mathbf{d^{\prime}}\rangle_{{}_{L_{2}}}}{||\mathbf{x}^{\prime}||||\mathbf{d}^{\prime}||},$
where $\mathbf{x}^{\prime},\mathbf{d}^{\prime}$ are derivative and $||\mathbf{x}||$ is the $L_{2}$ norm of $\mathbf{x}\in L_{2}([0,1])$. When multivariate functional data are considered, the inner product employed is the sum of the marginal inner product accross the dimension i.e. $\langle\mathbf{x},\mathbf{d}\rangle_{\mathcal{H}^{d}}=\sum_{i=1}^{d}\langle\mathbf{x}^{(i)},\mathbf{d}^{(i)}\rangle_{\mathcal{H}}$. The tree structure is equivalent to our algorithms and will be further detail in Section 3. For more detail about FIF specification, we refer the reader to Staerman et al., (2019).
2.2 The Signature Method
The signature of a path is a sequence of iterated integrals that captures important information about the path’s geometric and topological features (Lyons et al.,, 2007; Fermanian,, 2021).
Definition 2.1.
Let $\mathbf{X}\in\mathcal{F}^{d}$ be a r.v. of bounded variation. For any set of coordinates $\{i_{1},\ldots,i_{k}\}\subset\{1,\ldots,d\}^{k},k\in\mathbb{N}_{*}$, and $[s,t]\subset[0,1]$, the associated coordinate signature is defined by
$S_{(i_{1},\ldots,i_{k})}(\mathbf{X})_{[s,t]}$ =∫⋯∫$=\underset{s\leq u_{1}<\ldots<u_{k}\leq t}{\idotsint}\mathrm{d}X_{u_{1}}^{i_{1}}\ldots\mathrm{d}X_{u_{k}}^{i_{k}}.$
Furthermore, the signature of $\mathbf{X}$ is defined as the infinite collection of coordinate signature
$S(\mathbf{X})=\big{(}$ $1,S_{1}(\mathbf{X}),\ldots,S_{d}(\mathbf{X}),$
$S_{(1,1)}(\mathbf{X}),S_{(1,2)}(\mathbf{X}),\ldots,S_{(d,d)}(\mathbf{X}),$
$S_{(1,1,1)}(\mathbf{X}),S_{(1,2,1)}(\mathbf{X}),\ldots,S_{(d,d,d)}(\mathbf{X}),\ldots\big{)}.$
Truncated Signature. In practice, to be computable the truncated signature is used. Given an order of truncation $k\in\mathbb{N}_{*}$, it is defined as the signature vector of finite length of dimension $C=\left|\{1,\ldots,d\}^{k}\right|$ given by
$S^{k}(\mathbf{X})=\big{(}$ $1,S_{1}(\mathbf{X}),\ldots,S_{d}(\mathbf{X}),S_{(1,1)}(\mathbf{X}),S_{(1,2)}(\mathbf{X}),\ldots,$
$S_{(d,d)}(\mathbf{X}),\ldots,S_{\underbrace{(d,\ldots,d)}_{k}}(\mathbf{X})\big{)}\in\mathbb{R}^{C}.$
Kernel Truncated Signature. The signature can be seen as a feature map that embeds a function or a path into the tensor algebra. The truncated signature kernel has been recently introduced and studied in Király and Oberhauser, (2019). It is defined as the map $K^{k}:\mathcal{F}^{d}\times\mathcal{F}^{d}\rightarrow\mathbb{R}$ such that
$K^{k}(\mathbf{X},\mathbf{Y})=\langle S^{k}(\mathbf{X}),S^{k}(\mathbf{Y})\rangle.$ (1)
One may refer to Lee and Oberhauser, (2023) for an account of the untruncated kernel signature. See also Section 1 in the Appendix for further details about the signature and its properties.
3 Signature Isolation Forest Methods
With Kernel Signature Isolation Forest (K-SIF), we aim to leverage the truncated kernel signature (Király and Oberhauser,, 2019) to overcome the linearity constraint imposed by the inner product in FIF. In contrast to FIF, which explores only one function characteristic at each node using a unique inner product with the function sampled in the dictionary, K-SIF captures significantly more information. This is achieved by computing several coefficient signatures, summarizing multiple data attributes at each node.
The ability to explore non-linear directions in the feature space through the kernel signature makes the algorithm more efficient for an equivalent computational cost (refer to Figure 6 in the Appendix).
With Signature Isolation Forest (SIF), the aim is to leverage the coordinate signature and remove any restrictions imposed by FIF and K-SIF. We replace the use of a dictionary $\mathcal{D}$ and rely on the coordinate signature only to detect anomalies. The intuition is to provide a more straightforward, entirely data-driven solution and remove any a priori required choice.
The construction procedures of (K-)SIF are now described.
The objective of (K-)SIF is to construct a collection of (kernel) signature isolation trees, abbreviated by (k)si-trees, obtained from $\mathcal{X}=\{{\bf x}_{1},\dots,{\bf x}_{n}\}$, a training sample composed of independent realizations of ${\bf X}$ taking values in the space $\mathcal{F}^{d}$. If the dataset is too big, then a subsampling step can be performed to avoid masking or swamping effects (Chandola et al.,, 2009). This provides a subsample $\mathcal{X}_{\mathrm{sub}}=\{{\bf x}_{i_{1}},\dots,{\bf x}_{i_{m}}\}$ of size $m$ drawn randomly and uniformly to construct each tree. The global structure of a tree remains equivalent whether for a ksi-tree or a si-tree and is formally defined as follows.
Definition 3.1.
A (k)si-tree $\tau$ of depth $P\geq 1$ is a binary tree corresponding to a nested sequence of partitions of the functional space $\mathcal{F}^{d}$.
Tree Construction. Each (k)si-tree is obtained by recursively filtering $\mathcal{X}_{\mathrm{sub}}$, by means of a signature-based criterion, in a top-down fashion according to the procedure here presented. Such a tree starts with a root node, or initial node, corresponding to the entire feature space and given by $\mathcal{C}_{0,0}=\mathcal{F}^{d}$. Any other node encountered during the procedure will be denoted by the pair $(p,q)$ where $p$ is the index of the depth of the node (i.e. at which splitting node the algorithm has branched) with $0\leq p<P$, while $q$ is linked to the subset $\mathcal{C}_{p,q}\subset\mathcal{F}^{d}$ with $10\leq q\leq 2^{p}-$ 1. A non-terminal node $(p,q)$ has two children corresponding to two disjoint subset $\mathcal{C}_{p+1,2q}$ and $\mathcal{C}_{p+1,2q+1}$ such that $\mathcal{C}_{p,q}=\mathcal{C}_{p+1,2q}\cup\mathcal{C}_{p+1,2q+1}$. A node $(p,q)$ is said to be terminal if it has no children.
Splitting Criterion. The difference between a ksi-tree and a si-tree lies in the splitting criterion computed at each node of both trees. Below, we detail the splitting procedures occurring at a particular internal node for both algorithms.
Kernel Signature Isolation Forest. Due to the truncated signature use, the depth of this truncation level $k$ must be a priori chosen. Whilst the SIF does not require a dictionary specification, the K-SIF approaches algorithm involves as parameters a dictionary $\mathcal{D}\subset\mathcal{F}^{d}$, which is chosen to be rich enough to represent properties of the data considered and a probability measure $\boldsymbol{\nu}$ on $\mathcal{D}$. Here we use the three standard dictionaries of FIF (Staerman et al.,, 2019): ‘Brownian’, a dictionary of standard Brownian motion paths (corresponding to the space of continuous functions with the Wiener measure $\boldsymbol{\nu}$), ‘Gaussian wavelets’ representing a Mexican hat wavelet basis and ‘Cosine’ that is a cosine basis. In the latter two dictionary cases the probability measure $\boldsymbol{\nu}$ is usually uniform over the class of dictionary functions, which is suitable as an uniformative measure a priori.
The dataset at node $(p,q)$ is denoted at $\mathcal{X}_{p,q}$. At iteration $q+2^{p}$ of the ksi-tree construction, a split variable is selected by drawing a function $\mathbf{d}$ from $\mathcal{D}\subset\mathcal{F}^{d}$ according to a distribution $\boldsymbol{\nu}$. The data are then projected on such dictionary but through the signature kernel embedding, i.e. for a given ${\bf d}\in\mathcal{D}$ (where ${\bf d}$ represents a function sampled from $\mathcal{D}$ at each split) the projection of a sole function ${\bf x}$ considering the kernel signature is given according to the truncated signature kernel defined earlier in Eqn. 1 applied to $\mathbf{x}$ and dictionary item $\mathbf{d}$.This is performed $\forall{\bf x}\in\mathcal{X}_{p,q}$ and $k$ is a priori chosen.
The following step is to choose uniformly and at random a split value $\gamma$ such that
$\gamma\in\left[\min_{{\bf x}\in\mathcal{X}_{p,q}}\langle S^{k}({\bf x}),S^{k}({\bf d})\rangle,\;\;\max_{{\bf x}\in\mathcal{X}_{p,q}}\langle S^{k}({\bf x}),S^{k}({\bf d})\rangle\right].$
Then the algorithm can split $\mathcal{X}_{p,q}$ and generate a new node where the the children subsets are then defined by $\mathcal{C}_{p+1,2q}=\mathcal{C}_{p,q}\cap C_{\text{K-SIF}}^{L}$ and $\mathcal{C}_{p+1,2q+1}=\mathcal{C}_{p,q}\cap C_{\text{K-SIF}}^{R}$ with $L$ and $R$ indicating left and right subset splits given as
$\begin{split}C_{\text{K-SIF}}^{L}=&\{{\bf x}\in\mathcal{F}^{d}:\langle S^{k}({\bf x}),S^{k}({\bf d})\rangle\leq\gamma\}\\ C_{\text{K-SIF}}^{R}=&\{{\bf x}\in\mathcal{F}^{d}:\langle S^{k}({\bf x}),S^{k}({\bf d})\rangle>\gamma\}.\end{split}$
The children datasets are given as
$\begin{split}\mathcal{X}_{p+1,2q}&=\mathcal{X}_{p,q}\cap\;\mathcal{C}_{p+1,2q}\\ \mathcal{X}_{p+1,2q+1}&=\mathcal{X}_{p,q}\cap\;\mathcal{C}_{p+1,2q+1}.\end{split}$
A ksi-tree is built by iterating this procedure until all training data curves are isolated.
Signature Isolation Forest. In contrast to ksi-trees, the definition of a si-tree will rely on using a different split criterion, expressed according to the following procedure. Once again, the depth of the truncation level $k$ has to be chosen a priori, but no dictionaries nor distribution has to be set. Here, at each internal node $(p,q)$, a coordinate $(i_{1},\ldots,i_{\ell})$ is chosen randomly and uniformly in the set
$\{(i_{1},\ldots,i_{\ell})\in\llbracket 1,d\rrbracket^{\ell};\quad 1\leq\ell\leq k\}.$
Note that the coordinates may be chosen according to a specific law with additional a priori knowledge of the data, allowing the coordinate signature to discriminate particular aspects of the functions. Thereafter, this coordinate signature is computed across $\mathcal{X}_{p,q}$ and a split value $\gamma$ is chosen uniformly such that
$\gamma\in\left[\min_{{\bf x}\in\mathcal{X}_{p,q}}S_{(i_{1},\ldots,i_{\ell})}(\mathbf{x}),\;\;\max_{{\bf x}\in\mathcal{X}_{p,q}}S_{(i_{1},\ldots,i_{\ell})}(\mathbf{x})\right].$
Then the algorithm can split $\mathcal{X}_{p,q}$ and generate a new node where the the children subsets are then defined by $\mathcal{C}_{p+1,2q}=\mathcal{C}_{p,q}\cap C_{\text{SIF}}^{L}$ and $\mathcal{C}_{p+1,2q+1}=\mathcal{C}_{p,q}\cap C_{\text{SIF}}^{R}$ with $L$ and $R$ indicating left and right subset splits given as
$\begin{split}C_{\text{SIF}}^{L}=&\{{\bf x}\in\mathcal{F}^{d}:S_{(i_{1},\dots,i_{\ell})}({\bf x})\leq\gamma\}\\ C_{\text{SIF}}^{R}=&\{{\bf x}\in\mathcal{F}^{d}:S_{(i_{1},\dots,i_{\ell})}({\bf x})>\gamma\}.\end{split}$
The children datasets are given as
$\begin{split}\mathcal{X}_{p+1,2q}&=\mathcal{X}_{p,q}\cap\;\mathcal{C}_{p+1,2q}\\ \mathcal{X}_{p+1,2q+1}&=\mathcal{X}_{p,q}\cap\;\mathcal{C}_{p+1,2q+1}.\end{split}$
A si-tree is also built by iterating this procedure until all training data curves are isolated. The two algorithms of K-SIF and SIF are summarized in Section 1 in the Appendix.
Anomaly Score. As the terminal nodes of a (k)si-trees $\tau$ form a partition of the feature space, we can define the piecewise constant function $h_{\tau}:\mathcal{F}^{d}\rightarrow\mathbb{N}$ by:
$h_{\tau}({\bf x})=p\;\text{iff}\;{\bf x}\in\mathcal{C}_{p,q}\;\text{and}\;(p,q)\;\text{is a terminal node.}$.
This random path length offers an indication for the degree of abnormality: the more abnormal ${\bf x}$ is, the higher the probability that the quantity $h_{\tau}({\bf x})$ is small. K-SIF (resp. SIF) build a collection $\tau_{1},\ldots,\tau_{N}$ of $N$ ksi-trees (resp. si-tree). Given a $\mathbf{x}\in\mathcal{F}^{d}$, following Liu et al., (2008) and Staerman et al., (2019), we can define the monotone transformation of the averaged path length over the trees:
$s_{n}({\bf x})=2^{-\frac{1}{Nc(m)}\sum_{l=1}^{N}h_{\tau_{l}}({\bf x})},$
where $c(m)$ is the average path length of unsuccessful searches in a binary search tree and $m$ the size of the subsample linked to each tree.
Parameters of K-SIF and SIF. For both algorithms, key parameters typical of isolation forest-based methods, such as the number of trees $N$ or the subsample size $m$, must be pre-selected (Liu et al.,, 2008). The truncated level $k$ of the signature depth must also be chosen. In the case of K-SIF, similar to FIF, selecting a dictionary $\mathcal{D}$ and a distribution $\boldsymbol{\nu}$ is required. We opt to implement three standard dictionaries (also utilized in FIF): ‘Brownian’, representing a traditional Brownian motion path; ‘Cosine’, employing a cosine basis; and ‘Gaussian wavelets’, utilizing a Gaussian wavelets basis. In contrast, SIF does not require any of these sensitive parameters. We follow the approach outlined in Morrill et al., (2019) to enhance the algorithms’ performance by introducing a split window parameter $\omega$. This implies that, at each tree node, the truncated signature is computed on a randomly selected portion of the functions with a size of $\lfloor p/\omega\rfloor$.
4 Numerical Experiments
This section presents a series of numerical experiments supporting the proposed class of AD methods. We organize the experiments into three categories. First, a parameter sensitivity analysis sheds light on the behavior of the parameters algorithms and provides insights for achieving optimal performances. Subsequently, we compare (K-)SIF and FIF, illustrating the validity of such methods and showcasing the power of the signature in the context of AD solutions. Finally, we perform a benchmark that compares several existing AD methods on various real datasets. This benchmark aims to demonstrate the results in real applications and assess the overall performance of these algorithms relative to other AD methods documented in the literature.
4.1 Parameters Sensitivity Analysis
We investigate the behavior of K-SIF and SIF with respect to their two main parameters: the depth of the signature $k$ and the number of split windows $\omega$. For the sake of place, the experiment on the depth is postponed in Section 3.1 in the Appendix.
The Role of the Signature Split Window. The number of split windows allows the extraction of information over specific intervals (randomly selected) of the underlying data. Thus, at each tree node, the focus will be on a particular portion of the data, which is the same across all the sample curves for comparison purposes. This approach ensures that the analysis is performed on comparable sections of the data, providing a systematic way to examine and compare different intervals or features across the sample curves.
We explore the role of this parameter with two different datasets that reproduce two types of anomaly scenarios. The first considers isolated anomalies in a small interval, while the second contains persistent ones across all the function parametrization. In this way, we observe the behavior of K-SIF and SIF with respect to different types of anomalies.
The first dataset is constructed as follows. We simulate 100 constant functions. We then select at random 90% of these curves and Gaussian noise on a sub-interval; for the remaining 10% of the curves, we add Gaussian noise on another sub-interval, different from the first one. More precisely: 90% of the curves, considered as normal, are generated according to $\mathbf{x}(t)=b+\varepsilon(t)\mathbb{I}(t\in[0.3,0.6]),$ with $\varepsilon(t)\sim\mathcal{N}(0,1)$, $b\sim\mathcal{U}([0,100])$ and $\mathcal{U}$ representing the uniform distribution; 10% of the curves, considered as abnormal, are generated according to $\mathbf{x}(t)=b+\varepsilon(t)\mathbb{I}(t\in[0.7,0.8]),$ where $\varepsilon(t)\sim\mathcal{N}(0,1)$ and $b\sim\mathcal{U}([0,100])$.
The second dataset is constructed as follows. We generate 100 one-dimensional Brownian motion paths. We remark that a stochastic process $B_{t}$ follows a Brownian motion if it satisfies the following stochastic differential equation $dB_{t}=\mu dt+\sigma dX_{t},$ where $X_{t}$ is a Wiener process or Brownian motion (with distribution $\mathcal{N}(0,1)$).
We simulate at random 90% of the paths with $\mu=0$, $\sigma=0.5$, and consider them as normal data. Then, the remaining 10% are simulated with drift $\mu=0.2$, standard deviation $\sigma=0.4$, and considered abnormal data. We compute K-SIF with different numbers of split windows, varying from 1 to 10, with a truncation level set equal to 2 and $N=1,000$ the number of trees. The experiment is repeated $100$ times, and we report the averaged AUC under the ROC curves in Figure 1 for both datasets and three pre-selected dictionaries.
For the first dataset, where anomalies manifest in a small portion of the functions, increasing the number of splits significantly enhances the algorithm’s performance in detecting anomalies. The performance improvement shows a plateau after nine split windows. In the case of the second dataset with persistent anomalies, a higher number of split windows has a marginal impact on the algorithm’s performance, maintaining satisfactory results. Therefore, without prior knowledge about the data, opting for a relatively high number of split windows, such as 10, would ensure robust performance in both scenarios. Additionally, a more significant number of split windows enables the computation of the signature on a smaller portion of the functions, leading to improved computational efficiency.
Figure 1: AUC for the ROC curve w.r.t. the number of split window on the first (top) and the second (bottom) datasets for the three dictionaries.
4.2 (K-)SIF detects Swap Order Variation Changes
The signature method considers the order of occurring events in functional data. To investigate this phenomenon with our proposed class of algorithms, we define a synthetic dataset of $100$ smooth functions given by $\mathbf{x}(t)=30t^{q}(1-t)^{q},$ with $t\in[0,1]$ and $q$ equispaced in $[1,1.4]$. Then, we simulate the occurrences of events by adding Gaussian noise on different portions of the functions. We randomly select 90% of them and add Gaussian values on a sub-interval, i.e., $\mathbf{x}(t)=30t^{q}(1-t)^{q}+\varepsilon(t)\mathbb{I}(t\in[0.2,0.4]),$ where $\varepsilon(t)\sim\mathcal{N}(0,0.8)$. We consider the 10% remaining as abnormal by adding the same ‘events’ on another sub-interval compared to the first one, i.e., $\mathbf{x}(t)=30t^{q}(1-t)^{q}+\varepsilon(t)\mathbb{I}(t\in[0.6,0.8]),$ where $\varepsilon(t)\sim\mathcal{N}(0,0.8)$. See Figure 5 in the Appendix for a dataset visualization. To summarize, we have constructed two identical events occurring at different parts of the functions, leading to isolating anomalies. As presented in the introduction, this class of anomalous data are amongst the most challenging to identify.
We compute SIF, K-SIF and FIF with Brownian and Cosine dictionaries on these simulated datasets. As indicated in Section 4.1, for (K-)SIF, we choose $\omega=10$, the number of split windows, and $k=2$, the depth of the signature. In Figure 2, we report boxplots of the anomaly score returned by the algorithms for the normal data in purple and abnormal data in yellow. While this could appear a very simple task, in practice, it is highly challenging for an AD algorithm to differentiate such classes of curves. The introduction of the signature method should tackle precisely this type of scenario, since taking into account the order of the events.
Figure 2: Anomaly score for normal (purple) and abnormal (yellow) data for SIF, K-SIF and FIF with Brownian and Cosine dictionaries.
Table 1: AUROC of different anomaly detection methods calculated on the test set. Bold numbers correspond to the best result (higher is better).
Methods : SIF K-SIFGW FIFGW K-SIFC FIFC K-SIFB FIFB IF OCSVM fHD fSDO
Chinatown 1 0.90 0.84 0.99 0.9 1 0.83 0.81 0.6 0.73 1
Coffee 0.84 0.92 0.87 0.85 0.91 0.83 0.60 0.74 0.76 0.89 0.9
ECGFiveDays 0.93 0.90 0.88 0.92 0.95 0.90 0.93 0.85 0.93 0.84 0.96
ECG200 0.85 0.82 0.85 0.85 0.86 0.83 0.82 0.85 0.81 0.83 0.89
Handoutlines 0.84 0.83 0.83 0.83 0.83 0.82 0.8 0.84 0.83 0.82 0.83
SonyRobotAI1 0.99 0.96 0.98 0.95 0.93 0.95 0.76 0.96 0.96 0.88 0.96
SonyRobotAI2 0.93 0.89 0.88 0.92 0.80 0.93 0.84 0.87 0.84 0.84 0.88
StarLightCurves 0.80 0.75 0.71 0.76 0.67 0.76 0.77 0.72 0.71 0.71 0.79
TwoLeadECG 0.92 0.92 0.76 0.92 0.74 0.92 1 0.78 0.54 0.63 1
ECG5000 0.90 0.92 0.89 0.97 0.88 0.91 0.88 0.94 0.94 0.97 0.92
Using both dictionaries, FIF fails to detect this anomaly as it is not designed to handle these type of phenomena. In contrast, K-SIF and SIF produce significantly distinguished scores between normal and abnormal data, efficiently classifying the second data class as anomaly.
4.3 Real-data Anomaly Detection Benchmark
To evaluate the effectiveness of the proposed (K-)SIF algorithms and provide a comparison with FIF, we perform a comparative analysis using ten anomaly detection datasets constructed in Staerman et al., (2019) and sourced from the UCR repository (Chen et al.,, 2015). In contrast to Staerman et al., (2019), we do not use a training/test part since the labels are not used for the training and train and evaluate models on the training data only. We evaluate the algorithms’ performance by quantifying the Area under the ROC curves. For completeness, the FPR at 95% TPR and the Area under the PR curves are given in Tables 2 and 3 in the Appendix.
K-SIF is examined in the context of three introduced dictionaries: Brownian, Cosine and Gaussian wavelets, as done for FIF. The parameters are configured with $N=100$, $m=\text{min}(256,n)$, the height limit set to $\lceil\log_{2}(m)\rceil$ for both FIF and K-SIF/SIF. The number of split windows is fixed with $\omega=10$, and the depth is fixed with $k=3$. We compare the K-SIF/SIF methods against two widely used multivariate anomaly detection techniques and two functional depths with default settings. The multivariate methods, namely, Isolation Forest (IF; Liu et al.,, 2008) and one-class support vector machine (OCSVM; Schölkopf et al.,, 2001), are applied following dimension reduction via Functional PCA. We retain 20 principal components with the largest eigenvalues using the Haar basis. The depths considered include the random projection halfspace depth (Cuevas et al.,, 2007) and the functional Stahel-Donoho outlyingness (Hubert et al.,, 2015). Due to the relative small size of datasets, see Table 1 in the Appendix, deep learning based method performs poorly as shown in Tables 5, 6 and 7 in the Appendix.
On one hand, Figure 3 illustrates the performance disparity between FIF and K-SIF using the Brownian dictionary. Notably, K-SIF exhibits a significant performance advantage over FIF. This observation underscores the effectiveness of the signature kernel in improving FIF’s performance across most datasets, emphasizing the advantages of utilizing it over a simple inner product. On the other hand, considering the intricacy of functional data, no unique method is expected to outperform others universally.
Figure 3: Barplot of performance differences with AUC between K-SIF and FIF with a Brownian motion kernel (positive means K-SIF performs better), the inner product chose for FIF is L2 (top) and L2 of derivative (bottom).
However, SIF demonstrates strong performance in most cases, achieving the best results for five datasets. In contrast to FIF and K-SIF, it shows robustness to the variety of datasets while not being drastically affected by the choice of the parameters involved in FIF (dictionary and inner product) and K-SIF (dictionary).
5 Discussion & Conclusion
This work presents two novel anomaly detection algorithms, K-SIF and SIF, rooted in the isolation forest structure and the signature approach from rough path theory. Our contributions extend Functional Isolation Forest in two vital dimensions: incorporating non-linear properties in data for improved adaptability to challenging datasets and introducing an entirely data-driven technique free from predefined dictionaries. Such flexibility accommodates diverse data patterns and reduces the risk of overlooking certain types of anomalies. In this way, more complex real data pattern scenarios can be analysed, where non-linearity is highly present and, further, the unsupervised settings lacking labels highly affect the AD task. We demonstrate the advantages of utilizing K-SIF over FIF through a comprehensive parameter analysis, with consistent outperformance on real-world datasets. Notably, SIF achieves state-of-the-art performance while maintaining simplicity and computational efficiency, underscoring its effectiveness in functional anomaly detection. This work offers valuable advancements in anomaly detection methodologies, providing robust solutions for complex and diverse datasets.